# Diving Deep into Regions: Exploiting Regional Information Transformer for Single Image Deraining

### v1

@@ Abstract
Transformer-based Single Image Deraining (SID) methods have achieved remarkable success, primarily attributed to their robust capability in capturing long-range interactions. However, we’ve noticed that current methods handle rain-affected and unaffected regions concurrently, overlooking the disparities between these areas, resulting in confusion between rain streaks and background parts, and inabilities to obtain effective interactions, ultimately resulting in suboptimal deraining outcomes. To address the above issue, we introduce the Region Transformer (Regformer), a novel SID method that underlines the importance of independently processing rain-affected and unaffected regions while considering their combined impact for high-quality image reconstruction. The crux of our method is the innovative Region Transformer Block (RTB), which integrates a Region Masked Attention (RMA) mechanism and a Mixed Gate Forward Block (MGFB). Our RTB is used for attention selection of rain-affected and unaffected regions and local modeling of mixed scales. The RMA generates attention maps tailored to these two regions and their interactions, enabling our model to capture comprehensive features essential for rain removal. To better recover high-frequency textures and capture more local details, we develop the MGFB as a compensation module to complete local mixed scale modeling. Extensive experiments demonstrate that our model reaches state-of-the-art performance, significantly improving the image deraining quality. Our code and trained models will be publicly available at https://github.com/ztMotaLee/Regformer.
1 Introduction
@@
Single Image Deraining (SID) remains a critical and persistently challenging task in the field of computer vision, primarily due to its extensive applications in fields such as surveillance Dey & Bhattacharjee (2020), autonomous driving Sun et al. (2019), and remote sensing Li et al. (2020). The inclusion of rain in images not only diminishes visual quality but also exerts a substantial influence on the effectiveness of numerous high-level vision tasks Chen et al. (2023a); Zhou et al. (2023). Eliminating rain streaks from rainy images poses an inherent challenge owing to the ill-posed nature of this problem, intricate patterns, occlusions, and diverse densities of raindrops, all of which collectively contribute to the complexity of the deraining process.
@@
The emergence of deep learning has spearheaded significant progress in SID, with a multitude of Convolutional Neural Networks (CNN)-based methods being introduced Hu et al. (2019); Li et al. (2018); Yang et al. (2020). These methods have delivered promising outcomes across different scenarios. Beyond CNN-based approaches, Generative Adversarial Network (GAN)-based methods Qian et al. (2018); Zhang et al. (2019) have demonstrated their effectiveness in addressing the SID task. Through the use of adversarial loss, these models can produce visually appealing results, frequently surpassing traditional CNN-based techniques in preserving image details and structure. More recently, the remarkable success of transformer architecture in a broad range of computer vision tasks has prompted researchers to investigate its potential for SID Chen et al. (2023b); Jiang et al. (2022); Xiao et al. (2022).
Refer to caption Refer to captionRefer to caption(CVPR 22)(CVPR 23)(TPAMI 22)(Ours)(CVPR 22)(CVPR 23)(Ours)(TPAMI 22)
@@
Despite the substantial strides made with both transformer-based and CNN-based models in SID, when dealing with the complex and changeable real-world SID problem, especially when the rain streaks are very similar to the light spots in the image content, these methods often yield unsatisfactory results, as can be seen from Fig 2. Upon a closer examination, it can be observed that current models do not differentiate between the rain-affected and unaffected during processing. However, there are significant disparities between the unaffected regions of an image and the rain-affected areas, so it’s crucial to exploit the unique feature information inherent in these separate regions. Therefore, we believe that the distinct processing requirements of rain-affected and unaffected regions in images necessitate a more nuanced treatment than provided by current methods, as shown in Figure 1(a) and (b).
@@
Refer to caption rainy img_140 Wang et al. (2019) Refer to caption Refer to caption GT (PSNR/SSIM) RESCAN (25.17/0.9084) Li et al. (2018) Refer to caption Refer to caption Restormer (21.56/0.9041) Zamir et al. (2022) DualGCN (23.82/0.9291) Fu et al. (2021) Refer to caption Refer to caption SPDNet (23.51/0.9109) Yi et al. (2021) IDT (20.90/0.9125) Xiao et al. (2022) Refer to caption Refer to caption DRSformer (19.38/0.8894) Chen et al. (2023b) Ours (33.96/0.9682)
Figure 2: Visual comparison of different methods on SPA-Data Wang et al. (2019) dataset. Clearly, our method can better preserve the original unaffected details of the input image compared to other methods.
In addition, we introduce a novel module named Mixed Gate Forward Block (MGFB) to refine and enhance our model’s feature extraction capabilities. MGFB is expertly designed to process features across varying receptive fields. This technique models local interactions more effectively, thereby enabling the model to gain a comprehensive understanding of the image context by considering multiple-scale local features. With the embedded Mixed-Scale Modeling process, MGFB can better recover textures and preserve local details. The outcome is a highly refined feature set, resulting in more robust and detailed representations. In conjunction, the RTB with its RMA mechanism and the MGFB module empowers our Regformer model to overcome the constraints faced by existing methods, further leading to state-of-the-art performance, as shown in Figure 1(c).
In summary, the contributions of our work are three-fold.
@@
  • Our model achieves state-of-the-art performance on six public datasets, including four synthetic datasets and two real-world datasets. Our pretrained models and code will be publicly available.
2 Related Work
@@
2.1 Rain Streak and Raindrop Removal
@@
In recent years, deep learning-based approaches have outperformed their conventional counterparts in rain streak removal. These methods employ semi-supervised methods Wei et al. (2019), self-supervised module Huang et al. (2021), patch-based priors for both the unaffected and rain layers Huang et al. (2022), recurrent squeeze-and-excitation contextual dilated networks Li et al. (2018), conditional GAN-based models Zhang et al. (2019), and two-step rain removal techniques Ahn et al. (2022); Lin et al. (2020) to achieve promising results. As a unique rain degradation phenomenon, raindrop removal presents unique challenges that most rain streak removal algorithms cannot address directly. Consequently, researchers have proposed dedicated techniques for raindrop detection and removal, e.g., attentive GAN-based networks Qian et al. (2018); Zhang et al. (2019); Wei et al. (2021), CNN-based networks with shape-driven attention Quan et al. (2019), and two-stage video-based raindrop removal approaches Chen et al. (2022); Yan et al. (2022).
2.2 Vision Transformers in image deraining
@@
Following the Transformer network Vaswani et al. (2017), researchers have endeavored to replace CNNs with Transformers as the backbone structure for vision tasks. Dosovitskiy et al. Dosovitskiy et al. (2021) introduced the ViT by dividing an image into non-overlapping small patches and treating each patch as a token in the original transformer. Liu et al. Liu et al. (2021) proposed the Swin Transformer to further reduce the computational complexity of ViT. The Swin Transformer employs multi-headed self-attention in local windows rather than the global approach used in ViT. This modification reduces computation cost, while the shifted window approach compensates for any loss of global information.
@@
In the domain of image rain removal, several notable contributions have been made. For instance, Jiang et al. Jiang et al. (2022) proposed a dynamic-associated deraining network that integrates self-attention into a Transformer alongside an unaffected recovery network. Following this, Xiao et al. Xiao et al. (2022) introduced the Image Deraining Transformer (IDT), which employs both window-based and spatial-based dual Transformers, achieving remarkable results. Further advancements were made by Zamir et al. Zamir et al. (2022), who presented an efficient transformer model named Restormer. By incorporating innovative designs in the building blocks, Restormer can capture long-range pixel interactions while maintaining applicability to large images, which significantly contributes to subsequent image restoration. Building upon the foundation laid by Restormer, Chen et al. Chen et al. (2023b) put forward a sparse transformer. This model further explores useful self-attention values for feature aggregation and introduces a learnable top-k selection operator to procure more detailed features. Despite these efforts, current methodologies often stumble in comprehensively addressing the distinct regions within an image and their interactions, curtailing image deraining quality, as depicted in Figure 2. These limitations serve as the impetus for our Regformer model.
3 Method
@@
3.1 Overall Architecture
The overall pipeline of our Regformer model, as shown in Figure 3(a), employs an encoder-decoder architecture. Given an input image of size $H$$\times$$W$$\times$$C$, where $H$, $W$, and $C$ represent the height, width, and channel number of the image, respectively, we first preprocess the image using a 3$\times$3 convolution to extract the shallow feature. Note that we save a feature map during this stage, which is later utilized in the decoder part to obtain the mask. Subsequently, the feature map is passed through the RTB to extract global image features. Following each RTB in the encoder, we perform downsampling using 1$\times$1 convolutions and pixel-unshuffle operations, similar to methods Chen et al. (2023b); Wang et al. (2022); Xiao et al. (2022); Zamir et al. (2022). Concurrently, we introduce skip connections to bridge intermediate features and maintain network stability. Each RTB comprises one RMA, one MGFB, and two normalization layers. Note that we directly use a tensor with all 1 values as a mask at the encoder stage, since we have not yet accessed the restored feature maps.
In the decoder, we use $1\times 1$ convolution to reduce channels after processing the skip connections. Our Region Transformer Cascade (RTC) in the decoder consists of three RTBs and a $1\times 1$ convolution. Specifically, we duplicate the input feature maps and feed them into two RTBs separately. Within these two RTBs, we perform rain-affected and unaffected attention masking, resulting in two feature maps corresponding to the rain-affected and unaffected regions. A residual connection is then applied to avoid feature loss. We concatenate this resulting matrix and pass it through a $1\times 1$ convolution to reduce the channels. Finally, an RTB without a masking mechanism is used to integrate the three regional feature maps, yielding a comprehensive map containing rich feature information from the three distinct regions. The above process can be expressed as:
$F_{final}={RTB}({Conv}_{1\times 1}(Concat({RTB}_{U}(I),{RTB}_{R}(I),I))),$ (1)
@@ 3.1 Overall Architecture
3.2 Region Transformer Block
@@
3.2.1 Region Masked Attention
@@
Previous works did not fully consider the unique characteristics of different regions within an image. This oversight often resulted in a lack of differentiation between the rain-affected and unaffected areas, which in turn limited the effectiveness of their deraining outcomes. To address the limitation of previous works, we propose region masked attention, as illustrated in Figure 3(b). We first pass the input feature map through a 1x1 convolution to obtain the query (Q), key (K), and value (V) matrices. We then apply a 3 $\times$3 depth-wise separable convolution as a preprocessing step. After getting the processed result, we divide Q, K, and V to generate three independent projection matrices of key-value pairs. It is essential to use masking to Q and K after this step.
@@
where $I$ and $I^{\prime}$ denote the shallow features and the restored features in RTC, respectively; $R$ and $U$ denote the rain masks and unaffected region masks, respectively; $T$ signifies the application of dynamic thresholds; the $Binarize(\cdot)$ performs binarization, and ${Rep}_{C}(1)$ indicates the process of copying a single channel into C channels to align them. We use a dynamic threshold, to allow the network to adaptively distinguish the rain region and the unaffected region, and we binarize the segmentation result into a matrix of 0 or 1. In this case, the area where the value is set to 0 is not easily selected when processing the attention mechanism. However, this threshold cannot completely distinguish the rain region and unaffected region, which is shown in figure 6 and 7, so the RTB with the background mask is also needed. As a result, we derive the desired masks for rain-affected and unaffected regions.
Region Masked Attention. We then perform element-wise multiplication of the $Q$ and $K$ key-value pairs with the corresponding region mask to obtain the masked $Q$ and $K$ key-value pairs. Subsequently, we reshape $K$ to enable matrix multiplication, resulting in a region mask matrix with dimensions equal to the number of channels ($\hat{C}$). By performing matrix multiplication with this matrix and the $V$ key-value pair, we obtain the reconstructed attention map. The above process can be expressed as:
@@
3.2.2 Mixed Gate Forward Block
In the field of SID, one key point is how to conduct broader modeling for rain streaks of different scales. However, existing methods have not paid much attention to this Zamir et al. (2022), or are only limited to a specific scale modeling Chen et al. (2023b); Xiao et al. (2022), which will make the model perform poorly when retaining rain-like features, as can be seen from figure 2. In order to solve this problem, we design the Mixed Gate Forward Block (MGFB) with a Mixed Scale Modeling process to effectively perceive the local mixed-scale features in adjacent pixel areas, as depicted in Figure 3(c). Specifically, the input fused feature map is first passed through a 1$\times$1 convolution to expand the channels, and we use a 3$\times$3 depth-wise convolution to complete the preliminary feature extraction. This can be described as:
@@ 3.2.2 Mixed Gate Forward Block
$F=Activation(DWConv_{k_{1}\times k_{1}}(M)+M)\odot\prod_{i=2}^{n}(DWConv_{k_{i}\times k_{i}}(M)+M),$ (5)
@@
Refer to caption rainy img_465 Wang et al. (2019) Refer to caption Refer to caption GT (PSNR/SSIM) RESCAN (24.04/0.9024) Li et al. (2018) Refer to caption Refer to caption Restormer (23.48/0.9102) Zamir et al. (2022) DualGCN (22.09/0.8831) Fu et al. (2021) Refer to caption Refer to caption SPDNet (20.46/0.0.8593) Yi et al. (2021) IDT (27.60/0.9273) Xiao et al. (2022) Refer to caption Refer to caption DRSformer (23.79/0.9140) Chen et al. (2023b) Ours (38.26/0.9775)
Figure 5: Visual comparison on SPA-Data Wang et al. (2019) dataset. To facilitate the display of the results, we rotate the local graph 90 degrees clockwise. Unlike previous methods, our method successfully preserves the original details of the input image, compared to the GT.
4 Experiments
4.1 Experimental Settings
Evaluated Datasets. We evaluate the performance of our model on several public benchmarks, including Rain200L Yang et al. (2017), Rain200H Yang et al. (2017), DID-Data Zhang & Patel (2018), DDN-Data Fu et al. (2017) and SPA-Data Wang et al. (2019). Specifically, Rain200L and Rain200H consist of 1,800 synthetic rainy image pairs for training and 200 synthetic test image pairs. DID-Data and DDN-Data contain 12,000 and 12,600 synthetic rainy image pairs for training, respectively, along with 1,200 and 1,400 rainy image pairs for testing. Meanwhile, SPA-Data is a real-world dataset that includes 638,492 training image pairs and 1,000 testing image pairs. We further evaluate Regformer on a real-world raindrop dataset AGAN-Data Qian et al. (2018), which includes 1,110 image pairs.
Compared Methods. For rain streak removal, we compare our Regformer model with a variety of state-of-the-art methods, including prior knowledge-based approaches like DSC Luo et al. (2015) and GMM Li et al. (2016), CNN-based methods such as DDN Fu et al. (2017), RESCAN Li et al. (2018), PReNet Ren et al. (2019), MSPFN Jiang et al. (2020), RCDNet Wang et al. (2020), MPRNet Zamir et al. (2021), DualGCN Fu et al. (2021) and SPDNet Yi et al. (2021). In addition, we consider Transformer-based methods, including Uformer Wang et al. (2022), Restormer Zamir et al. (2022), IDT Xiao et al. (2022) and DRSformer Chen et al. (2023b). We refer to the experimental data provided by DRSformer and adhere to their original experimental settings for further analysis. Our experimental results are presented in Table 1. In addition, we also test our model on one real-world raindrop dataset AGAN-Data Qian et al. (2018) to measure our model performance more broadly. We compare our model with five methods, including Eigen et al. Eigen et al. (2013), Pix2Pix Isola et al. (2017), AttentGAN Qian et al. (2018), Quan et al. Quan et al. (2019) and IDT Xiao et al. (2022). Our experimental results are presented in Table 2. We employ PSNR (Peak Signal-to-Noise Ratio) Q. & M. (2008) and SSIM (Structural Similarity Index Measure) Wang et al. (2004) as the primary evaluation metrics. Note that we calculate the metrics on the Y channel of YCbCr like the previous works Chen et al. (2023b); Wang et al. (2020); Yang et al. (2017).
Implementation and Training Details. In our Regformer model, we configure the layers L0, L1, L2 and L3 to 4, 4, 4, 4, and the corresponding attention heads to 6, 6, 6, 6. Additionally, we set the initial channel to 48 for Regformer. During training, we employ the AdamW Loshchilov & Hutter (2019) optimizer with parameters $\beta 1$ = 0.9, $\beta 2$ = 0.999, and weight decay = 1e-4. In the experiment part, we set $n$ = 2, $k_{1}$ = 3 and $k_{2}$ = 5 in the MGFB. Please note that in different tasks or different datasets, we may modify these hyperparameters to achieve better results. Concurrently, we follow DRSformer’s settings Chen et al. (2023b), which utilize the L1 loss as the loss function. We set the initial learning rate to 3e-4 and apply a cosine annealing algorithm to gradually decay it to 1e-6. The patch size is set to $128\times 128$, and we perform a total of 300K iterations during training using two NVIDIA GeForce RTX 3090 GPUs with 24GB memory each. Any settings that are not explicitly mentioned in this paper keep consistent with those used in DRSformer Chen et al. (2023b).
4.2 Deraining Performance Evaluations

### v2

# Exploiting Regional Information Transformer for Single Image Deraining

@@ Abstract
Transformer-based Single Image Deraining (SID) methods have achieved remarkable success, primarily attributed to their robust capability in capturing long-range interactions. However, we’ve noticed that current methods handle rain-affected and unaffected regions concurrently, overlooking the disparities between these areas, resulting in confusion between rain streaks and background parts, and inabilities to obtain effective interactions, ultimately resulting in suboptimal deraining outcomes. To address the above issue, we introduce the Region Transformer (Regformer), a novel SID method that underlines the importance of independently processing rain-affected and unaffected regions while considering their combined impact for high-quality image reconstruction. The crux of our method is the innovative Region Transformer Block (RTB), which integrates a Region Masked Attention (RMA) mechanism and a Mixed Gate Forward Block (MGFB). Our RTB is used for attention selection of rain-affected and unaffected regions and local modeling of mixed scales. The RMA generates attention maps tailored to these two regions and their interactions, enabling our model to capture comprehensive features essential for rain removal. To better recover high-frequency textures and capture more local details, we develop the MGFB as a compensation module to complete local mixed scale modeling. Extensive experiments demonstrate that our model reaches state-of-the-art performance, significantly improving the image deraining quality. Our code and trained models are publicly available at https://github.com/ztMotaLee/Regformer.
Index Terms:
Single Image Deraining, Vision Transformer, Region Attention.
I Introduction
@@
Single Image Deraining (SID) remains a critical and persistently challenging task in the field of computer vision, primarily due to its extensive applications in fields such as surveillance [1, 2], autonomous driving [3, 4, 5], and remote sensing [6, 7]. The inclusion of rain in images not only diminishes visual quality but also exerts a substantial influence on the effectiveness of numerous high-level vision tasks [8, 9]. Eliminating rain streaks from rainy images poses an inherent challenge owing to the ill-posed nature of this problem, intricate patterns, occlusions, and diverse densities of raindrops, all of which collectively contribute to the complexity of the deraining process.
@@
The emergence of deep learning has spearheaded significant progress in SID, with a multitude of Convolutional Neural Networks (CNN)-based methods being introduced [10, 11, 12]. These methods have delivered promising outcomes across different scenarios. Beyond CNN-based approaches, Generative Adversarial Network (GAN)-based methods [13, 14] have demonstrated their effectiveness in addressing the SID task. Through the use of adversarial loss, these models can produce visually appealing results, frequently surpassing traditional CNN-based techniques in preserving image details and structure. More recently, the remarkable success of transformer architecture in a broad range of computer vision tasks has prompted researchers to investigate its potential for SID [15, 16, 17].
Refer to caption Refer to captionRefer to caption
@@
Despite the substantial strides made with both transformer-based and CNN-based models in SID, when dealing with the complex and changeable real-world SID [18, 19] problem, especially when the rain streaks are very similar to the light spots in the image content, these methods often yield unsatisfactory results, as can be seen from Figures 4 and 5. Upon a closer examination, it can be observed that current models do not differentiate between the rain-affected and unaffected during processing. However, there are significant disparities between the unaffected regions of an image and the rain-affected areas, so it’s crucial to exploit the unique feature information inherent in these separate regions. Therefore, we believe that the distinct processing requirements of rain-affected and unaffected regions in images necessitate a more nuanced treatment than provided by current methods, as shown in Figure 1(a) and (b).
@@
In addition, we introduce a novel module named Mixed Gate Forward Block (MGFB) to refine and enhance our model’s feature extraction capabilities. MGFB is expertly designed to process features across varying receptive fields. This technique models local interactions more effectively, thereby enabling the model to gain a comprehensive understanding of the image context by considering multiple-scale local features. With the embedded Mixed-Scale Modeling process, MGFB can better recover textures and preserve local details. The outcome is a highly refined feature set, resulting in more robust and detailed representations. In conjunction, the RTB with its RMA mechanism and the MGFB module empowers our Regformer model to overcome the constraints faced by existing methods, further leading to state-of-the-art performance, as shown in Figure 1(c). In summary, the contributions of our work are three-fold:
@@
  • Our model achieves state-of-the-art performance on various synthetic and real-world datasets. Our pretrained models and code are made publicly available.
II Related Work
@@
II-A Rain Streak and Raindrop Removal
@@
In recent years, deep learning-based approaches have outperformed their conventional counterparts in rain streak removal. These methods employ semi-supervised methods [20], self-supervised module [21], patch-based priors for both the unaffected and rain layers [22], recurrent squeeze-and-excitation contextual dilated networks [11], conditional GAN-based models [14], and two-step rain removal techniques [23, 24] to achieve promising results. As a unique rain degradation phenomenon, raindrop removal presents unique challenges that most rain streak removal algorithms cannot address directly. Consequently, researchers have proposed dedicated techniques for raindrop detection and removal, e.g., attentive GAN-based networks [13, 14, 25], CNN-based networks with shape-driven attention [26], and two-stage video-based raindrop removal approaches [27, 28].
II-B Vision Transformers in Image Deraining
@@
Following the Transformer network [29], researchers have endeavored to replace CNNs with Transformers as the backbone structure for vision tasks [30, 31, 32]. Dosovitskiy et al. [33] introduced the ViT by dividing an image into non-overlapping small patches and treating each patch as a token in the original transformer. Liu et al. [34] proposed the Swin Transformer to further reduce the computational complexity of ViT. The Swin Transformer employs multi-headed self-attention in local windows rather than the global approach used in ViT. This modification reduces computation cost, while the shifted window approach compensates for any loss of global information.
@@
In the domain of image rain removal, several notable contributions have been made [35, 36, 37]. For instance, Jiang et al. [16] proposed a dynamic-associated deraining network that integrates self-attention into a Transformer alongside an unaffected recovery network. Following this, Xiao et al. [17] introduced the Image Deraining Transformer (IDT), which employs both window-based and spatial-based dual Transformers, achieving remarkable results. Further advancements were made by Zamir et al. [38], who presented an efficient transformer model named Restormer. By incorporating innovative designs in the building blocks, Restormer can capture long-range pixel interactions while maintaining applicability to large images, which significantly contributes to subsequent image restoration. Building upon the foundation laid by Restormer, Chen et al. [15] put forward a sparse transformer. This model further explores useful self-attention values for feature aggregation and introduces a learnable top-k selection operator to procure more detailed features. Despite these efforts, current methodologies often stumble in comprehensively addressing the distinct regions within an image and their interactions, curtailing image deraining quality, as depicted in Figures 4 and 5. These limitations serve as the impetus for our Regformer model.
III Method
@@
III-A Overall Architecture
The overall pipeline of our Regformer model, as shown in Figure 2(a), employs an encoder-decoder architecture. Given an input image of size $H$$\times$$W$$\times$$C$, where $H$, $W$, and $C$ represent the height, width, and channel number of the image, respectively, we first preprocess the image using a 3$\times$3 convolution to extract the shallow feature. Note that we save a feature map during this stage, which is later utilized in the decoder part to obtain the mask. Subsequently, the feature map is passed through the RTB to extract global image features.
Following each RTB in the encoder, we perform downsampling using 1$\times$1 convolutions and pixel-unshuffle operations, similar to methods [15, 39, 17, 38]. Concurrently, we introduce skip connections to bridge intermediate features and maintain network stability. Each RTB comprises one RMA, one MGFB, and two normalization layers. Note that we directly use a tensor with all 1 values as a mask at the encoder stage, since we have not yet accessed the restored feature maps.
In the decoder, we use $1\times 1$ convolution to reduce channels after processing the skip connections. Our Region Transformer Cascade (RTC) in decoder contains three RTBs and a $1\times 1$ convolution. Specifically, we duplicate the input feature maps and feed them into two RTBs separately. Within these two RTBs, we perform rain-affected and unaffected attention masking, resulting in two feature maps corresponding to the rain-affected and unaffected regions. A residual connection is then applied to avoid feature loss. We concatenate this resulting matrix and pass it through a $1\times 1$ convolution to reduce the channels. Finally, an RTB without a masking mechanism is used to integrate the three regional feature maps, yielding a comprehensive map containing rich feature information from the three distinct regions. The above process can be expressed as
$\footnotesize F_{final}={RTB}({Conv}_{1\times 1}(Concat({RTB}_{U}(I),{RTB}_{R}(I),I))),\vspace{+1mm}$ (1)
@@ III-A Overall Architecture
III-B Region Transformer Block
@@
III-B1 Region Masked Attention
Previous works did not fully consider the unique
@@ III-B1 Region Masked Attention
characteristics of different regions within an image. This oversight often resulted in a lack of differentiation between the rain-affected and unaffected areas, which in turn limited the effectiveness of their deraining outcomes. To address the limitation of previous works, we propose region masked attention, as illustrated in Figure 2(b). We first pass the input feature map through a 1x1 convolution to obtain the query (Q), key (K), and value (V) matrices. We then apply a 3 $\times$3 depth-wise separable convolution as a preprocessing step. After getting the processed result, we divide Q, K, and V to generate three independent projection matrices of key-value pairs. It is essential to use masking to Q and K after this step.
@@
where $I$ and $I^{\prime}$ denote the shallow features and the restored features in RTC, respectively; $R$ and $U$ denote the rain masks and unaffected region masks, respectively; $T$ signifies the application of dynamic thresholds; the $Binarize(\cdot)$ performs binarization, and ${Rep}_{C}(1)$ indicates the process of copying a single channel into C channels to align them. We use a dynamic threshold, to allow the network to adaptively distinguish the rain region and the unaffected region, and we binarize the segmentation result into a matrix of 0 or 1. In this case, the area where the value is set to 0 is not easily selected when processing the attention mechanism. However, this threshold cannot completely distinguish the rain region and unaffected region, which is shown in figure 9, so the RTB with the background mask is also needed. As a result, we derive the desired masks for rain-affected and unaffected regions.
Attention Mechanism. We then perform element-wise multiplication of the $Q$ and $K$ key-value pairs with the corresponding region mask to obtain the masked $Q$ and $K$ key-value pairs. Subsequently, we reshape $K$ to enable matrix multiplication, resulting in a region mask matrix with dimensions equal to the number of channels ($\hat{C}$). By performing matrix multiplication with this matrix and the $V$ key-value pair, we obtain the reconstructed attention map. The above process can be expressed as:
@@
III-B2 Mixed Gate Forward Block
In the field of SID, one key point is how to conduct broader modeling for rain streaks of different scales. However, existing methods have not paid much attention to this [38], or are only limited to a specific scale modeling [15, 17], which will make the model perform poorly when retaining rain streak-like content features and eliminating content-like rain streak noise, as can be seen from Figures 4 and 5. In order to solve this problem, we design the Mixed Gate Forward Block (MGFB) with a Mixed Scale Modeling process to effectively perceive the local mixed-scale features in adjacent pixel areas, as depicted in Figure 2(c). Specifically, the input fused feature map is first passed through a 1$\times$1 convolution to expand the channels, and we use a 3$\times$3 depth-wise convolution to complete the preliminary feature extraction. This can be described as:
@@ III-B2 Mixed Gate Forward Block
$F=Activation(DWConv_{k_{1}\times k_{1}}(M)+M)$ (5)
$\odot\prod_{i=2}^{n}(DWConv_{k_{i}\times k_{i}}(M)+M),$